# Large Language Models are Algorithmically Blind

### v1

@@ Abstract
Sohan Venkatesh1thanks: Correspondence: soh.venkatesh@gmail.com   Ashish Mahendran Kurapath1   Tejas Melkote1 1School of Computer Engineering, Manipal Institute of Technology, Bengaluru, India
@@ 3.1 Ground Truth Computation
$\text{Precision}=\frac{|\hat{E}\cap E^{*}|}{|\hat{E}|},\qquad\text{Recall}=\frac{|\hat{E}\cap E^{*}|}{|E^{*}|}$ (1)
$\text{F1}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (2)
@@
$\text{SHD}(\hat{G},G^{*})$ $=|\text{missing edges}|+|\text{extra edges}|$ (3)
$\quad+|\text{reversed edges}|$
@@ 3.3 Aggregation and Coverage Computation
$\text{Coverage}(M,D,A,m)$ $=\mathbf{1}\Big[\hat{\mu}_{D,A,m}\in$ (4)
$\quad\left[\hat{l}_{M,D,A,m},\,\hat{u}_{M,D,A,m}\right]\Big]$
@@ 4.1 Systematic Calibration Failure
Across all 1,664 comparisons, frontier LLMs achieve a mean calibrated coverage of 15.9%, meaning predicted ranges contain the true algorithmic mean fewer than 1 in 6 times. This represents an 84.1% failure rate. Table 1 reports per-model results.
@@ 4.4 Benchmark versus Synthetic Degradation
Figure 4: Coverage collapse across synthetic network sizes.
@@ 4.5 Memorization Probes
Figure 5: Predicted range width: benchmark vs. synthetic. Mean compression ratio 2.26× across all models.
@@ Appendix B Prompt Formulations and CV% analysis
Figure 7 shows the three prompt formulations used across all experimental conditions.
@@
We evaluate prompt robustness by querying each model with three distinct formulations per experimental condition and computing the coefficient of variation (CV%) of predicted range midpoints and widths across formulations. The three formulations vary along two dimensions: (1) specificity of metric naming, where f1 uses direct metric names, f2 adds algorithm intuition and context and f3 uses alternative uncertainty-focused phrasing; and (2) framing of the prediction task, ranging from direct numerical elicitation to open-ended range description.
@@ Appendix C LiNGAM Synthetic Failure Analysis
A per-model breakdown of LiNGAM calibrated coverage on synthetic data reveals a near-total collapse across all models.
Table 10: LiNGAM synthetic coverage breakdown.
(a) Per-Model LiNGAM Calibrated Coverage on Synthetic Data
@@
(b) Average Synthetic Coverage by Algorithm Algorithm Synthetic NOTEARS 21.9% FCI 10.9% PC 10.2% LiNGAM 3.9%     (c) Benchmark vs Synthetic Coverage Alg. Bench Synth Diff LiNGAM 27.1% 3.9% $23.2\%$ NOTEARS 20.1% 21.9% $1.8\%$ PC 12.2% 10.2% $2.0\%$ FCI 11.5% 10.9% $0.6\%$
Seven of eight models score 0% on LiNGAM synthetic data across all network sizes, as shown in Table 10(a). Claude is the sole exception, achieving partial coverage on recall and F1 while failing entirely on precision and SHD; this is addressed in Section 4.1 as algorithm-specific pattern matching rather than genuine understanding.
Table 10(b) ranks all four algorithms by average synthetic coverage. LiNGAM ranks last, performing 5.6 times worse than NOTEARS despite being evaluated on the same datasets and models. This gap cannot be attributed to dataset difficulty since all algorithms face identical conditions.
[…]
@@ Appendix D Algorithm versus LLM comparison
Two patterns are visible in Table 12. First, three of the four synthetic datasets rank among the four lowest positions, aligning with the benchmark/synthetic degradation reported in Section 4.4. Second, within benchmark datasets, coverage correlates inversely with network size: Asia (8 nodes, 23.4%) and Cancer (5 nodes, 21.9%) are the highest-performing benchmarks, while Hepar2 (70 nodes, 10.9%) is the lowest. Synthetic-12 is the sole exception, achieving 20.3% coverage comparable to small benchmarks, consistent with its network size falling within the range of well-documented benchmark graphs. The monotonic synthetic collapse from 20.3% to 6.2% across 12 to 60 nodes is reported separately in Section 4.4 and Figure 4.
@@ Appendix E Algorithmic Ground Truth and Statistical Analysis
Table 13: Mean F1 Score by Algorithm (100 runs $\times$ 13 datasets)
@@
(a) Mean Precision, Recall and F1 by algorithm aggregated across all datasets. FCI and PC achieve the highest recall, while LiNGAM and NOTEARS show lower recall with differing precision profiles reflecting their distinct assumptions.
@@
Figure 9: Structural Hamming Distance (mean SHD with 95% CIs) by algorithm. NOTEARS has the lowest mean SHD, followed by PC, FCI and LiNGAM, with LiNGAM exhibiting the highest variance.
@@
Figure 10: Precision, Recall and F1 by dataset and algorithm. Subplots show per-dataset variation across all four algorithms, highlighting dataset dependence and the role of graph complexity.
Figure 11: F1 score heatmap. NOTEARS dominates synthetic datasets, LiNGAM peaks on Survey and PC/FCI show consistent benchmark performance, supporting the memorization hypothesis.

### v2

@@ Abstract
Sohan Venkatesh1thanks: Correspondence: soh.venkatesh@gmail.com   Ashish Mahendran Kurapath1   Tejas Melkote1
1School of Computer Engineering, Manipal Institute of Technology, Bengaluru, India
@@ 3.1 Ground Truth Computation
$\text{Precision}=\frac{|\hat{E}\cap E^{*}|}{|\hat{E}|},\quad\text{Recall}=\frac{|\hat{E}\cap E^{*}|}{|E^{*}|},\quad\text{F1}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (1)
@@
$\text{SHD}(\hat{G},G^{*})=|\text{missing edges}|+|\text{extra edges}|+|\text{reversed edges}|$ (2)
@@ 3.3 Aggregation and Coverage Computation
$\text{Coverage}(M,D,A,m)=\mathbf{1}\!\left[\hat{\mu}_{D,A,m}\in\left[\hat{l}_{M,D,A,m},\,\hat{u}_{M,D,A,m}\right]\right]$ (3)
@@ 4.1 Systematic Calibration Failure
[…]
@@ 4.4 Benchmark versus Synthetic Degradation
Figure 4: Coverage collapse across synthetic networks.
@@ 4.5 Memorization Probes
Figure 5: Predicted range width: benchmark vs. synthetic datasets
[…]
@@ Appendix B Prompt Formulations and CV% analysis
We evaluate prompt robustness by querying each model with three distinct formulations per experimental condition and computing the coefficient of variation (CV%) of predicted range midpoints and widths across formulations. The three formulations vary along two dimensions: (1) specificity of metric naming, where f1 uses direct metric names, f2 adds algorithm intuition and context and f3 uses alternative uncertainty-focused phrasing; and (2) framing of the prediction task, ranging from direct numerical elicitation to open-ended range description. Figure 7 shows the three prompt formulations used across all experimental conditions.
@@ Appendix C LiNGAM Synthetic Failure Analysis
A per-model breakdown of LiNGAM calibrated coverage on synthetic data reveals a near-total collapse across all models. Seven of eight models score 0% on LiNGAM synthetic data across all network sizes, as shown in Table 10. Claude is the sole exception, achieving partial coverage on recall and F1 while failing entirely on precision and SHD; this is addressed in Section 4.1 as algorithm-specific pattern matching rather than genuine understanding.
Table 10: Per-Model LiNGAM Calibrated Coverage on Synthetic Data
@@
Tables 11(a) and (b) rank algorithms by average synthetic coverage and show the benchmark vs synthetic flip. LiNGAM ranks last, performing 5.6 times worse than NOTEARS despite being evaluated on the same datasets and models. This gap cannot be attributed to dataset difficulty since all algorithms face identical conditions. LiNGAM holds the highest benchmark coverage of any algorithm (27.1%) yet the lowest synthetic coverage (3.9%), a 23.2 percentage point inversion that is the largest dataset type gap across all algorithm model combinations. Every other algorithm shows near-stable or modest decline. The inversion suggests that LLMs may be retrieving memorized statistics and fail completely when that scaffold is removed. The $7\times$ drop from benchmark to synthetic quantifies the scale of LLMs’ dependence on training data exposure rather than algorithmic understanding.
Tables 11(a) and (b) show algorithm-level synthetic coverage and the benchmark vs synthetic flip. LiNGAM ranks last at 3.9%, performing 5.6 times worse than NOTEARS despite identical evaluation conditions, yet holds the highest benchmark coverage (27.1%), a 23.2 percentage point inversion that is the largest dataset-type gap across all algorithm-model combinations. Every other algorithm shows near-stable or modest decline, ruling out a general synthetic-data difficulty effect. The inversion suggests LLMs retrieve memorized benchmark statistics and fail completely when that scaffold is removed. The $7\times$ drop quantifies LLMs’ dependence on training data exposure rather than algorithmic understanding.
Table 11: Synthetic coverage by algorithm (a) and benchmark vs synthetic comparison (b).
Algorithm Synthetic (%)
NOTEARS 21.9
FCI 10.9
PC 10.2
LiNGAM 3.9
Algorithm Benchmark (%) Synthetic (%) Diff.
LiNGAM 27.1 3.9 23.2-23.2
NOTEARS 20.1 21.9 +1.8+1.8
PC 12.2 10.2 2.0-2.0
FCI 11.5 10.9 0.6-0.6
@@ Appendix D Algorithm versus LLM comparison
Two patterns are visible in Table 13. First, three of the four synthetic datasets rank among the four lowest positions, aligning with the benchmark/synthetic degradation reported in Section 4.4. Second, within benchmark datasets, coverage correlates inversely with network size: Asia (8 nodes, 23.4%) and Cancer (5 nodes, 21.9%) are the highest-performing benchmarks, while Hepar2 (70 nodes, 10.9%) is the lowest. Synthetic-12 is the sole exception, achieving 20.3% coverage comparable to small benchmarks, consistent with its network size falling within the range of well-documented benchmark graphs. The monotonic synthetic collapse from 20.3% to 6.2% across 12 to 60 nodes is reported separately in Section 4.4.
@@ Appendix E Algorithmic Ground Truth and Statistical Analysis
Table 14: Mean F1 Score by Algorithm
@@
(a) Mean Precision, Recall and F1 by algorithm aggregated across all datasets. FCI and PC achieve the highest recall, while LiNGAM and NOTEARS show lower recall with differing precision profiles.
@@
(c) Structural Hamming Distance (mean SHD with 95% CIs) by algorithm. NOTEARS has the lowest mean SHD, followed by PC, FCI and LiNGAM, with LiNGAM exhibiting the highest variance.